what is the maximum amount of SF4 that can be made in the reaction 16.87g of NaF and 98.42g of SCI2

1
Expert's answer
2022-05-19T21:12:03-0400

3 SCl2 + 4 NaF = SF4 + S2Cl2 + 4 NaCl

1. We find the amount of NaF substance. for this we use the formula

n = m / M = 16.87 / 42 = 0.4 mol NaF.

Theoretically we find how many moles SF4 are formed. 

According to the reaction,

1 mole of SF4 is formed from 4 moles of NaF. How many moles of SF4 are formed from 0.4 moles of NaF? :

4mol NaF —> 1mol SF4

0.4mol NaF –> X

X= 0.4 * 1 / 4 = 0.1 mol SF4.

2. We find the amount of SCl2 substance. for this we use the formula

n = m / M = 98.42 / 103 = 0.96 mol SCl2. Theoretically we find how many moles SF4 are formed. 

According to the reaction,

1 mole of SF4 is formed from 3 moles of SCl2. How many moles of SF4 are formed from 0.96 moles of SCl2 ? :

3mol SCl2 —> 1mol SF4

0.96mol SCl2 –> X

X= 0.96 * 1 / 3 = 0.32 mol SF4.

3. In practice, we get the answer in which calculation the least SF4 is formed. that is, we obtain SF4 (0.1mol) calculated on NaF. because NaF is a limiting reagent in this reaction. (All of the NaF reacts. Some of the SCl2 increases.)

4. We find the mass of 0.1 mol of SF4. for this we use the formula m = n * M:

m = n * M(SF4) = 0.1 * 108 = 10.8 gr

ANSWER: 0.1mol or 10.8gr SF4